A rectangular painting with a width of x centimeters has an area x^2 + 50x square centimeters. Find a binomial that represents the length.
Answer by stanbon(75887)

A rectangular painting with a width of x centimeters has an area x^2 + 50x square centimeters. Find a binomial that represents the length.
------------------
Area of a rectagle = width*length
Your Problem:
x^2+50x = x*length
length = (x^2+50x)/x
length = x+50
=================
Cheers,
Stan H.